Which numbers are greater than 5 × 10 - 6?
Choose all answers that are correct.
A. 9 × 10 - 8
B. 7 × 10 - 6
C. 5 × 10 - 4
D. 4 × 10 - 6
E. 8 × 10 - 9

Found 2 solutions by purpleswimmergirl, Jc0110:
Answer by purpleswimmergirl(1)

First, we must solve 5 x 10 - 6.
(5x10) - 6 = 50 - 6 = 44
Now, we can work through A, B, C, D, and E
A.
(9x10) - 8 = 90 - 8 = 82
B.
(7x10) - 6 = 70 - 6 = 64
C.
(5x10) - 4 = 50 - 4 = 46
D.
(4x10) - 6 = 40 - 6 = 34
E.
(8x10) - 9 = 80 - 9 = 71
Therefore, A, B, C, and E are greater than 5 x 10 - 6.

Hope that helped.
Kind regards,
M

Answer by Jc0110(165)

Strictly obey BODMAS rules.
From question: 5 × 10 - 6 = 50 - 6 = 44
--------------------------------------------------------
From A.: 9 × 10 - 8 = 90 - 8 = 82
82>44
A. > Question
----------------------------------
From B.: 7 × 10 - 6 = 70 - 6 = 64
64>44
B. > Question
----------------------------------
From C.: 5 × 10 - 4 = 50 - 4 = 46
46>44
C. > Question
----------------------------------
From D.: 4 × 10 - 6 = 40 - 6 = 34
34<44
D. < Question
----------------------------------
From E.: 8 × 10 - 9 = 80 - 9 = 71
71>44
E. > Question
--------------------------------------------------------
Therefore, A., B., C. and E. are the answers.